# Global well-posedness for the generalized intermediate NLS with a nonvanishing condition at infinity

### v1

@@ 4. Global well-posedness
Our goal in this section is to establish Theorem 1.2. To do so, we derive the conservation laws for (gINLS) in the $\mathcal{Z}^{2}_{\rho}$ framework, which in particular are meaningful for solutions with nonvanishing boundary conditions at infinity. We then show that, for arbitrary real parameters $\alpha\neq 0$, $\beta\geq 0$, the Cauchy problem (gINLS) is globally well-posed in $\mathcal{Z}^{2}_{\rho}$. In addition, when $\alpha=\beta=1$ we obtain a uniform bound for global solutions to (INLS).
@@
Let $\delta>0$, $\alpha,\beta\in\mathbb{R}$, and $\rho>0$. Let $\phi\in\mathcal{Z}^{2}_{\rho}$, and let $u\in C([0,T_{\max});\mathcal{Z}^{2}_{\rho})$ be the unique maximal-lifespan solution to the Cauchy problem for (gINLS) with the initial condition $u(0)=\phi$ constructed in Theorem 1.1. Then,
@@ 2. $\mathrm{(ii)}$
    (4.1) $\mathcal{H}_{2}(u(t))\lesssim_{\alpha,\beta,\delta}\mathcal{H}_{2}(u_{0})+\int_{0}^{t}\Big(\mathcal{E}^{1}_{\rho}(u(s))+\big[\mathcal{E}^{1}_{\rho}(u(s))\big]^{\frac{3}{2}}\Big)\Big(1+\|\partial_{x}^{2}u(s)\|_{L^{2}}^{2}\Big)\,ds.$
@@ 3. $\mathrm{(iii)}$
    Assume $\alpha\neq 0$ and $\beta\geq 0$. Then, the solution $u$ obeys the following a priori bounds,
    (4.2) $\mathcal{E}^{1}_{\rho}(u(t))$ $\leq C_{1},$
@@
    for $t\in[0,T_{\max})$, where the constants $C_{1},C_{2},C_{3}>0$ depend only on $\alpha,\beta,\delta,\rho$ and $\mathcal{E}^{1}_{\rho}(\phi)$. In particular, $T_{\max}=\infty$.
To derive $\mathcal{H}_{1}(u)$ and $\mathcal{H}_{2}(u),$ we proceed by constructing modified energies through a cancellation method [shatah, colliander-kdv]. We begin with the time derivative of a basic quantity (either $\int|\partial_{x}u|^{2}dx$ or $\int|\partial_{x}^{2}u|^{2}dx$) and systematically introduce correctors to cancel out the problematic nonlinear terms. For $\mathcal{H}_{1}(u)$, this process is continued to provide a conserved quantity for general $\alpha,\beta$. Although this is not the case for $\mathcal{H}_{2}(u)$ in general, in the integrable case we can also derive an exact conserved quantity as in the following corollary.
@@
Let $\delta>0$, $\rho>0$ and $\phi\in\mathcal{Z}^{2}_{\rho}$. If $\beta=\pm|\alpha|$, the following quantity
@@
is well-defined in $\mathcal{Z}^{2}_{\rho}$ and gives a conservation law for (gINLS). Furthermore, if $\beta=|\alpha|>0$, the global solution $u\in C(\mathbb{R};\mathcal{Z}^{2}_{\rho})$ of (gINLS) satisfies
(4.4) $\sup_{t\in\mathbb{R}}\mathcal{E}^{2}_{\rho}(u(t))<\infty.$
Before proving Theorem 4.1, we collect some useful lemmas.
@@
Let $u\in C([0,T];\mathcal{Z}^{\infty}_{\rho})$ be a smooth solution of (gINLS). Then,
@@
$\lesssim\Big(\mathcal{E}^{1}_{\rho}(u)+\big[\mathcal{E}^{1}_{\rho}(u)\big]^{\frac{3}{2}}\Big)\Big(1+\|\partial_{x}^{2}u\|_{L^{2}}^{2}\Big).$
@@
$\partial_{t}|\mathcal{H}_{2}(u)|\lesssim_{\alpha,\beta}\sum_{j=1}^{5}|R_{j}(u)|\lesssim_{\alpha,\beta,\delta}\Big(\mathcal{E}^{1}_{\rho}(u)+\big[\mathcal{E}^{1}_{\rho}(u)\big]^{\frac{3}{2}}\Big)\Big(1+\|\partial_{x}^{2}u\|_{L^{2}}^{2}\Big).$
@@
$\lesssim\big[\mathcal{E}^{1}_{\rho}(\phi)\big]^{\frac{1}{2}}+\big[\mathcal{E}^{1}_{\rho}(\phi)\big]^{\frac{3}{2}}.$
@@
where $C_{0}>0$ and other implicit constants may depend on $\alpha,\beta,\delta,\rho,\mathcal{E}^{1}_{\rho}(\phi)$. In particular, we have
@@
Here, the implicit constants depend on $\alpha,\delta,\rho$ and $\mathcal{E}^{1}_{\rho}(u)$. As we did for (4.7), by Young’s inequality we deduce
$\Big|\mathcal{H}_{2}^{\mathrm{INLS}}(u(t))-\int|\partial_{x}^{2}u(t)|^{2}\Big|\leq\frac{1}{2}\int|\partial_{x}^{2}u(t)|^{2}+C_{4}\big(\alpha,\delta,\rho,\mathcal{E}^{1}_{\rho}(u(t))\big)\qquad(t\in[0,T])$
for some constant $C_{4}>0$. Then, from the a priori bound (4.2) for $\mathcal{E}^{1}_{\rho}(u(t))$ proven in Theorem 4.1 (here we use the defocusing assumption $\beta=|\alpha|>0$) and conservation of $\mathcal{H}_{2}^{\mathrm{INLS}}(u(t))$, we infer that
$\|\partial_{x}^{2}u(t)\|_{L^{2}}^{2}$ $\leq 2\mathcal{H}_{2}^{\mathrm{INLS}}(u(t))+2C_{4}\big(\alpha,\delta,\rho,\mathcal{E}^{1}_{\rho}(u(t))\big)$
$\leq 2\mathcal{H}_{2}^{\mathrm{INLS}}(\phi)+2C_{4}\big(\alpha,\delta,\rho,C_{1}(\alpha,\delta,\rho,\mathcal{E}^{1}_{\rho}(\phi)\big)$
$\leq 3\|\partial_{x}^{2}\phi\|_{L^{2}}^{2}+2C_{4}\big(\alpha,\delta,\rho,\mathcal{E}^{1}_{\rho}(\phi)\big)+2C_{4}\big(\alpha,\delta,\rho,C_{1}(\alpha,\delta,\rho,\mathcal{E}^{1}_{\rho}(\phi))\big),$
@@ Appendix A Fourier transform of the hyperbolic cotangent function
(i) The first inequality follows immediately from the definition of $c_{j}[f]$. We also see that $c_{j\pm 1}[f]\leq 2^{\varepsilon}c_{j}[f]$, which implies the second inequality. For the last one, we easily see $\mathcal{E}^{2}_{\rho}(f)\lesssim\|\mathbf{c}[f]\|_{\ell^{2}}^{2}$, while the converse is shown by the (discrete) Young inequality.
@@
$|c_{j}[f]-c_{j}[g]|\leq\sum_{k=0}^{\infty}2^{-\varepsilon|j-k|}\big\|P_{k}\partial_{x}(f-g)\big\|_{H^{1}}+2^{-\varepsilon j}\Big(\|f-g\|_{L^{\infty}}+\big\||f|^{2}-|g|^{2}\big\|_{L^{2}}^{\frac{1}{2}}\Big).$
@@
$\|P_{\leq j}f\|_{L^{\infty}}^{2}+\big\||P_{\leq j}f|^{2}-\rho^{2}\big\|_{L^{2}}$ $\leq\|P_{\leq j}f\|_{L^{\infty}}^{2}+\big\||f|^{2}-\rho^{2}\big\|_{L^{2}}+\big(\|P_{\leq j}f\|_{L^{\infty}}+\|f\|_{L^{\infty}}\big)\|P_{\geq j+1}f\|_{L^{2}}$
$\lesssim\|f\|_{L^{\infty}}^{2}+\big\||f|^{2}-\rho^{2}\big\|_{L^{2}}+\sum_{k=j+1}^{\infty}\|P_{k}f\|_{L^{2}}^{2}$
@@
By Bernstein’s inequality, the $\mathcal{Z}^{3}_{\rho}$ bound from (3.24), and Lemma 3.4 (iii), we see that
$\|P_{k}\partial_{x}^{2}(u^{j}-u^{j-1})\|_{L^{\infty}_{T_{*}}L^{2}}$ $\lesssim 2^{-k}\big(\|\partial_{x}u^{j}\|_{L^{\infty}_{T_{*}}H^{2}}+\|\partial_{x}u^{j-1}\|_{L^{\infty}_{T_{*}}H^{2}}\big)\lesssim 2^{-k}\big(\mathcal{E}^{3}_{\rho}(P_{\leq j}\phi)+\mathcal{E}^{3}_{\rho}(P_{\leq j-1}\phi)\big)^{\frac{1}{2}}$
@@
$\|P_{k}\partial_{x}^{2}(u^{j}-u^{j-1})\|_{L^{\infty}_{T}L^{2}}$ $\lesssim 2^{k}\|\partial_{x}(u^{j}-u^{j-1})\|_{L^{\infty}_{T}L^{2}}\lesssim_{M_{0}}2^{k}d^{1}_{\rho}(P_{\leq j}\phi,P_{\leq j-1}\phi)$

### v2

@@ 3. $\mathrm{(iii)}$
We conclude this section by showing the convergence result.
Proof of Corollary 1.3.
Let $\{\phi_{\delta}\}_{\delta\in[1,\infty]}$, $\{u_{\delta}\}_{\delta\in[1,\infty]}$ be as in the statement of the corollary, and take $M_{0}>0$ such that $\mathcal{E}^{2}(\phi_{\infty})\leq M_{0}^{2}$. Then, there exists $\delta_{*}\in[1,\infty)$ such that $\mathcal{E}^{2}(\phi_{\delta})\leq 2M_{0}^{2}$ and $\mathcal{E}^{2}(P_{\leq j}\phi_{\delta})\leq 6M_{0}^{2}$ for any $\delta\in[\delta_{*},\infty]$ and $j\in\mathbb{Z}_{\geq 0}$. Let $u_{\delta}^{j}$ denote the (unique) solution to (gINLS) ($\delta\in[1,\infty)$) or (gCM) ($\delta=\infty$) with $u_{\delta}^{j}(0)=P_{\leq j}\phi_{\delta}$. From the proof of Theorem 1.1, there exists $T_{*}\gtrsim(1+M_{0}^{4})^{-1}$ such that for any $\delta\in[\delta_{*},\infty]$ and $j\in\mathbb{Z}_{\geq 0}$, these solutions exist on $0,T_{*}]$ (in particular $T_{\delta,\max}>T_{*}$) and satisfy
(3.32) $\|\mathcal{E}^{2}(u_{\delta})\|_{L^{\infty}_{T_{*}}}+\|\mathcal{E}^{2}(u_{\delta}^{j})\|_{L^{\infty}_{T_{*}}}$ $\lesssim M_{0}^{2},$
(3.33) $\|\mathcal{E}^{3}(u_{\delta}^{j})\|_{L^{\infty}_{T_{*}}}$ $\lesssim\mathcal{E}^{3}(P_{\leq j}\phi_{\delta})\lesssim 2^{2j}M_{0}^{2},$
(3.34) $\|d^{1}(u_{\delta}^{j},u_{\delta})\|_{L^{\infty}_{T_{*}}}$ $\lesssim_{M_{0}}d^{1}(P_{\leq j}\phi_{\delta},\phi_{\delta}),$
(3.35) $\|\partial_{x}^{2}(u_{\delta}-u_{\delta}^{j})\|_{L^{\infty}_{T_{*}}L^{2}}$ $\lesssim_{M_{0}}\|\mathbf{c}[\phi_{\delta}]\|_{\ell^{2}(\mathbb{Z}_{>j})}.$
Indeed, these estimates have been shown in (3.24), (3.25), (3.27), and (3.29), respectively, with the constants independent of $\delta\in[1,\infty]$.
We next prove the difference estimate
(3.36) $\|d^{1}(u_{\delta},u_{\infty})\|_{L^{\infty}_{T_{*}}}\lesssim(1+M_{0})d^{1}(\phi_{\delta},\phi_{\infty})+\delta^{-\frac{1}{2}}(1+M_{0}),\qquad\delta\in[\delta_{*},\infty).$
To see (3.36), it suffices to prove
(3.37) $\|d^{1}(u_{\delta}^{j},u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}}\lesssim(1+M_{0})d^{1}(P_{\leq j}\phi_{\delta},P_{\leq j}\phi_{\infty})+\delta^{-\frac{1}{2}}(1+M_{0}),\qquad\delta\in[\delta_{*},\infty),\ j\geq 0$
and take the limit $j\to\infty$ with the estimate (3.34). Then, since $u_{\delta}^{j}$ is smooth, we can follow the calculations in the proof of Proposition 3.3. In fact, the main part of the proof involving the modified energy estimate remains unchanged; we only have to replace the difference of $\mathcal{N}_{\delta}$’s for two solutions by a single $\mathcal{N}_{\delta}(u_{\delta}^{j})$. From Lemma 2.1 and (3.32) we have
$\|\mathcal{N}_{\delta}(u_{\delta}^{j})\|_{L^{\infty}_{T_{*}}H^{2}}\lesssim\delta^{-1}M_{0}^{3}.$
Then, the previous argument for Proposition 3.3 implies the estimates
$\|d^{1}(u_{\delta}^{j},u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}}$ $\lesssim(1+M_{0})d^{1}(P_{\leq j}\phi_{\delta},P_{\leq j}\phi_{\infty})+\delta^{-1}T_{*}(1+M_{0})M_{0}^{3}+\|\partial_{x}(u_{\delta}^{j}-u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}L^{2}},$
$\|\partial_{x}(u_{\delta}^{j}-u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}L^{2}}^{2}$ $\lesssim\|\partial_{x}P_{\leq j}(\phi_{\delta}-\phi_{\infty})\|_{L^{2}}^{2}+\delta^{-1}T_{*}(M_{0}^{4}+M_{0}^{6})$
$\quad+\big(T_{*}(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})+2^{-n_{2}}M_{0}^{2}\big)\|d^{1}(u_{\delta}^{j},u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}}^{2}$
instead of (3.18) and (3.23), respectively. Choosing $n_{2}$ appropriately and modifying $T_{*}\sim(1+M_{0}^{4})^{-1}$ if necessary, we obtain (3.37).
In view of (3.36), the convergence $\|d^{2}(u_{\delta},u_{\infty})\|_{L^{\infty}_{T_{*}}}\to 0$ follows once $\|\partial_{x}^{2}(u_{\delta}-u_{\infty})\|_{L^{\infty}_{T_{*}}L^{2}}\to 0$ is verified. To prove the latter convergence, we follow the argument showing continuous dependence in Step 3 of the proof of Theorem 1.1. Using interpolation, (3.33), and (3.37) in place of (3.27), we have the fixed-$j$ difference estimate corresponding to (3.31):
$\|\partial_{x}^{2}(u_{\delta}^{j}-u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}L^{2}}^{2}$ $\leq\big(\|\partial_{x}^{3}u_{\delta}^{j}\|_{L^{\infty}_{T_{*}}L^{2}}+\|\partial_{x}^{3}u_{\infty}^{j}\|_{L^{\infty}_{T_{*}}L^{2}}\big)\|d^{1}(u_{\delta}^{j},u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}}$
$\lesssim_{M_{0}}2^{j}\big(d^{1}(P_{\leq j}\phi_{\delta},P_{\leq j}\phi_{\infty})+\delta^{-\frac{1}{2}}\big)$
$\lesssim_{M_{0}}2^{j}\big(d^{1}(\phi_{\delta},\phi_{\infty})+\delta^{-\frac{1}{2}}\big).$
Combining it with (3.35), we obtain
$\limsup_{\delta\to\infty}\|\partial_{x}^{2}(u_{\delta}-u_{\infty})\|_{L^{\infty}_{T_{*}}L^{2}}$ $\leq\limsup_{\delta\to\infty}\Big(\|\partial_{x}^{2}(u_{\delta}-u_{\delta}^{j})\|_{L^{\infty}_{T*}L^{2}}+\|\partial_{x}^{2}(u_{\infty}-u_{\infty}^{j})\|_{L^{\infty}_{T*}L^{2}}+\|\partial_{x}^{2}(u_{\delta}^{j}-u_{\infty}^{j})\|_{L^{\infty}_{T_{*}}L^{2}}\Big)$
$\lesssim_{M_{0}}\limsup_{\delta\to\infty}\Big(\|\mathbf{c}[\phi_{\delta}]\|_{\ell^{2}(\mathbb{Z}_{>j})}+\|\mathbf{c}[\phi_{\infty}]\|_{\ell^{2}(\mathbb{Z}_{>j})}+\big[2^{j}\big(d^{1}(\phi_{\delta},\phi_{\infty})+\delta^{-\frac{1}{2}}\big)\big]^{\frac{1}{2}}\Big)$
$\lesssim\|\mathbf{c}[\phi_{\infty}]\|_{\ell^{2}(\mathbb{Z}_{>j})}\ \to 0\qquad(j\to\infty).$
We have thus proven that $\lim\limits_{\delta\to\infty}\|d^{2}(u_{\delta},u_{\infty})\|_{L^{\infty}_{T_{*}}}=0$.
Now, we take $T\in(0,T_{\infty,\max})$ arbitrarily. Since $\mathcal{E}^{2}(u_{\infty}(t))$ is bounded on $[0,T]$, repeating the above procedure finitely many times we see that there exists $\delta_{**}\in[1,\infty)$ such that $T_{\delta,\max}>T$ for any $\delta\in[\delta_{**},\infty]$ and $\lim\limits_{\delta\to\infty}\|d^{2}(u_{\delta},u_{\infty})\|_{L^{\infty}_{T}}=0$. This completes the proof of Corollary 1.3. ∎
@@ 4. Global well-posedness
Our goal in this section is to establish Theorem 1.4. To do so, we derive the conservation laws for (gINLS) in the $\mathcal{Z}^{2}$ framework; in particular, they are meaningful for solutions with nonvanishing boundary conditions at infinity. We then show that, for any real parameters $\alpha\neq 0$, $\beta\geq 0$, the Cauchy problem for (gINLS) is globally well-posed in $\mathcal{Z}^{2}_{*}$. In addition, when $\alpha=\beta=1$ we obtain a uniform bound for global solutions to (INLS). Our argument also yields the corresponding results for (gCM) and (CM-defoc).
@@
Let $\delta>0$, $\alpha,\beta\in\mathbb{R}$. Let $\phi\in\mathcal{Z}^{2}_{\rho}$ for some $\rho\geq 0$, and let $u\in C([0,T_{\max});\mathcal{Z}^{2}_{\rho})$ be the unique maximal solution to the Cauchy problem for (gINLS) with the initial condition $u(0)=\phi$ constructed in Theorem 1.1.
@@ 2. $\mathrm{(ii)}$
    (4.1) $\mathcal{H}_{2}(u(t))\lesssim_{\alpha,\beta,\delta}\mathcal{H}_{2}(u_{0})+\int_{0}^{t}\Big(\mathcal{E}^{1}(u(s))+\big[\mathcal{E}^{1}(u(s))\big]^{\frac{3}{2}}\Big)\Big(1+\|\partial_{x}^{2}u(s)\|_{L^{2}}^{2}\Big)\,ds.$
@@ 3. $\mathrm{(iii)}$
    Assume $\alpha\neq 0$, $\beta\geq 0$ and $\rho>0$. Then, the solution $u$ obeys the following a priori bounds:
    (4.2) $\mathcal{E}^{1}(u(t))$ $\leq C_{1},$
@@
    for $t\in[0,T_{\max})$, where the constants $C_{1},C_{2},C_{3}>0$ depend only on $\alpha,\beta,\delta,\rho$ and $\mathcal{E}^{1}(\phi)$. In particular, $T_{\max}=\infty$.
Furthermore, for (gCM) we have the same results as above, with $T_{\delta}$ replaced by $H$ and with the constants being independent of $\delta$.
To derive $\mathcal{H}_{1}(u)$ and $\mathcal{H}_{2}(u),$ we proceed by constructing modified energies through a cancellation method [40, 13]. We begin with the time derivative of the leading term (either $\int|\partial_{x}u|^{2}dx$ or $\int|\partial_{x}^{2}u|^{2}dx$) and systematically introduce correctors to cancel out the problematic nonlinear terms. For $\mathcal{H}_{1}(u)$, this process is continued to provide a conserved quantity for general $\alpha,\beta$. Although this is not the case for $\mathcal{H}_{2}(u)$ in general, in the integrable case we can also derive an exact conserved quantity as in the following corollary.
@@
Let $\delta>0$ and $\phi\in\mathcal{Z}^{2}_{\rho}$ for some $\rho\geq 0$. If $\beta=\pm|\alpha|$, the quantity
@@
is well-defined in $\mathcal{Z}^{2}_{\rho}$ and is a conserved quantity for (gINLS). If $\beta=|\alpha|>0$ and $\rho>0$, the global solution $u\in C(\mathbb{R};\mathcal{Z}^{2}_{\rho})$ to (gINLS) satisfies
(4.4) $\sup_{t\in\mathbb{R}}\mathcal{E}^{2}(u(t))<\infty.$
Furthermore, for (gCM) we have the same results as above, with $T_{\delta}$ replaced by $H$.
Remark 4.1.
For vanishing solutions (in $H^{s}$), infinitely many conservation laws for (CM-defoc), (CM-foc) and for (INLS) have been obtained in [25, 19] and in [9], respectively. In particular, a conserved quantity at the $H^{k}(\mathbb{R})$ level is given by $\|\mathcal{L}_{u(t)}^{k}u(t)\|_{L^{2}}^{2}$, where $\mathcal{L}_{u}$ is a suitable Lax operator (see (1.5) for the case of (INLS)). In [11, Lemma 3.5], Chen found conserved quantities at the $\mathcal{Z}^{1}$ and $\mathcal{Z}^{2}$ levels for (CM-defoc), which were obtained by expanding $\|\mathcal{L}_{u}u\|_{L^{2}}$, $\|\mathcal{L}_{u}^{2}u\|_{L^{2}}$ and “renormalizing” them for nonvanishing solutions, i.e., replacing $\Pi(|u|^{2})$ with $\Pi(|u|^{2}-1)+1$, where $\Pi=\frac{1}{2}(1+iH)$ is the Cauchy-Szegő projection and is a bounded operator on $L^{2}(\mathbb{R})$. It is then natural to ask whether conservation laws for nonvanishing solutions to (INLS) can be obtained by similar renormalization procedure. However, for (INLS) the operator $\frac{1}{2}(1+iT_{\delta})$ corresponding to $\Pi$ becomes unbounded on $L^{2}(\mathbb{R})$, and there is no apparent way of renormalizing $\|\mathcal{L}_{u}^{k}u\|_{L^{2}}$ for $u\in\mathcal{Z}^{k}_{\rho}$. Instead, in Theorem 4.1 and Corollary 4.2, we construct conserved quantities directly by using the idea of modified energy method.
From now on, we only consider (gINLS); however, we can apply exactly the same argument to (gCM) by replacing $T_{\delta}$ with $H$. Before proving Theorem 4.1, we collect some useful lemmas.
@@
Let $u\in C([0,T];\mathcal{Z}^{\infty})$ be a smooth solution of (gINLS). Then,
@@
$\lesssim\Big(\mathcal{E}^{1}(u)+\big[\mathcal{E}^{1}(u)\big]^{\frac{3}{2}}\Big)\Big(1+\|\partial_{x}^{2}u\|_{L^{2}}^{2}\Big).$
@@
$\partial_{t}|\mathcal{H}_{2}(u)|\lesssim_{\alpha,\beta}\sum_{j=1}^{5}|R_{j}(u)|\lesssim_{\alpha,\beta,\delta}\Big(\mathcal{E}^{1}(u)+\big[\mathcal{E}^{1}(u)\big]^{\frac{3}{2}}\Big)\Big(1+\|\partial_{x}^{2}u\|_{L^{2}}^{2}\Big).$
@@
$\lesssim\big[\mathcal{E}^{1}(\phi)\big]^{\frac{1}{2}}+\big[\mathcal{E}^{1}(\phi)\big]^{\frac{3}{2}}.$
@@
where $C_{0}>0$ and other implicit constants may depend on $\alpha,\beta,\delta,\rho,\mathcal{E}^{1}(\phi)$. In particular, we have
@@
Here, the implicit constants depend on $\alpha,\delta,\rho$ and $\mathcal{E}^{1}(u)$. As we did for (4.7), by Young’s inequality we deduce
$\Big|\mathcal{H}_{2}^{\mathrm{INLS}}(u(t))-\int|\partial_{x}^{2}u(t)|^{2}\Big|\leq\frac{1}{2}\int|\partial_{x}^{2}u(t)|^{2}+C_{4}\big(\alpha,\delta,\rho,\mathcal{E}^{1}(u(t))\big)\qquad(t\in[0,T])$
for some constant $C_{4}>0$. Then, from the a priori bound (4.2) for $\mathcal{E}^{1}(u(t))$ proven in Theorem 4.1 (here we use the defocusing assumption $\beta=|\alpha|>0$) and conservation of $\mathcal{H}_{2}^{\mathrm{INLS}}(u(t))$, we infer that
$\|\partial_{x}^{2}u(t)\|_{L^{2}}^{2}$ $\leq 2\mathcal{H}_{2}^{\mathrm{INLS}}(u(t))+2C_{4}\big(\alpha,\delta,\rho,\mathcal{E}^{1}(u(t))\big)$
$\leq 2\mathcal{H}_{2}^{\mathrm{INLS}}(\phi)+2C_{4}\big(\alpha,\delta,\rho,C_{1}(\alpha,\delta,\rho,\mathcal{E}^{1}(\phi)\big)$
$\leq 3\|\partial_{x}^{2}\phi\|_{L^{2}}^{2}+2C_{4}\big(\alpha,\delta,\rho,\mathcal{E}^{1}(\phi)\big)+2C_{4}\big(\alpha,\delta,\rho,C_{1}(\alpha,\delta,\rho,\mathcal{E}^{1}(\phi))\big),$
@@ Appendix A Fourier transform of the hyperbolic cotangent function
(i) The first inequality follows immediately from the definition of $c_{j}[f]$. We also see that $c_{j\pm 1}[f]\leq 2^{\varepsilon}c_{j}[f]$, which implies the second inequality. For the last one, we easily see $\mathcal{E}^{2}(f)\lesssim\|\mathbf{c}[f]\|_{\ell^{2}}^{2}$, while the converse is shown by the (discrete) Young inequality.
@@
$|c_{j}[f]-c_{j}[g]|\leq\sum_{k=0}^{\infty}2^{-\varepsilon|j-k|}\big\|P_{k}\partial_{x}(f-g)\big\|_{H^{1}}+2^{-\varepsilon j}\Big(\|f-g\|_{L^{\infty}}+\big\|(|f|^{2}-\rho_{f}^{2})-(|g|^{2}-\rho_{g}^{2})\big\|_{L^{2}}^{\frac{1}{2}}\Big).$
@@
$\|P_{\leq j}f\|_{L^{\infty}}^{2}+\big\||P_{\leq j}f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}$ $\leq\|P_{\leq j}f\|_{L^{\infty}}^{2}+\big\||f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}+\big(\|P_{\leq j}f\|_{L^{\infty}}+\|f\|_{L^{\infty}}\big)\|P_{\geq j+1}f\|_{L^{2}}$
$\lesssim\|f\|_{L^{\infty}}^{2}+\big\||f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}+\sum_{k=j+1}^{\infty}\|P_{k}f\|_{L^{2}}^{2}$
@@
By Bernstein’s inequality, the $\mathcal{Z}^{3}$ bound from (3.25), and Lemma 3.4 (iii), we see that
$\|P_{k}\partial_{x}^{2}(u^{j}-u^{j-1})\|_{L^{\infty}_{T_{*}}L^{2}}$ $\lesssim 2^{-k}\big(\|\partial_{x}u^{j}\|_{L^{\infty}_{T_{*}}H^{2}}+\|\partial_{x}u^{j-1}\|_{L^{\infty}_{T_{*}}H^{2}}\big)\lesssim 2^{-k}\big(\mathcal{E}^{3}(P_{\leq j}\phi)+\mathcal{E}^{3}(P_{\leq j-1}\phi)\big)^{\frac{1}{2}}$
@@
$\|P_{k}\partial_{x}^{2}(u^{j}-u^{j-1})\|_{L^{\infty}_{T}L^{2}}$ $\lesssim 2^{k}\|\partial_{x}(u^{j}-u^{j-1})\|_{L^{\infty}_{T}L^{2}}\lesssim_{M_{0}}2^{k}d^{1}(P_{\leq j}\phi,P_{\leq j-1}\phi)$